OpenStudy (anonymous):

Hello - can you help with the final answer please? I am not sure if it should be 168m x 168m = 336m or 168m^2 Formula: F^2 = I^2 - 2ad F=velocity speed I = initial velocity -2ad = 2 x deceleration x distanced travelled F^2 = 13.89^2 - 2 x 6 x 2.08 = ? I hope you understand this. Thank you so much

OpenStudy (dumbcow):

Are you trying to find final velocity, F ? just put it in calculator

OpenStudy (littlebird):

Just remember you have to find the square root of both sides as well

OpenStudy (dumbcow):

oh i see where you got stuck ... yeah take sqrt of 168 F^2 = 168 F = sqrt(168)

OpenStudy (anonymous):

is that 28224?

OpenStudy (dumbcow):

no thats 168^2 what number squared = 168 examples: sqrt(25) = 5 sqrt(36) = 6 sqrt(81) = 9

OpenStudy (anonymous):

so answer is 12.96??

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

Thank you so much for your help!!!

OpenStudy (dumbcow):

yw :)